OpenStudy (anonymous):

what is an equation of a line passing through (3,3) and (4,7) ?

OpenStudy (goformit100):

First of all for solving this question do you know the conception of : Basic Straight Lines

OpenStudy (anonymous):

m=4

OpenStudy (texaschic101):

m = (y2 - y1)/(x2-x1)......set 1(3,3)...set 2(4,7) m = (7 - 3)/(4 - 3) m = 4/1 = 4 y = mx + b.....(3,3) 3 = 4(3) + b 3 = 12 +b 3 - 12 = b -9 = b y = 4x - 9 <-- answer